OpenStudy (anonymous):

In a right triangle, the square of the hypothenuse is equal to the sum of the squares of the legs. If the legs of a right triangle are 6 and 8, find the lenght of the hypothenuse

OpenStudy (cgreenwade2000):

The square of the hyp^2 = legA^2 + legB^2

OpenStudy (cgreenwade2000):

You may have seen it before as a^2 + b^2 = c^2 where a and b are legs of a right triangle and c is the hypotenuse.

OpenStudy (anonymous):

I got 10 is that right?

OpenStudy (cgreenwade2000):

Yes. 6^2 + 8^2 = c^2 100= c^2 10=c

OpenStudy (anonymous):

Thank you for your help.